# PCL-Indexability and Whittle Index for Restless Bandits with General Observation Models

### v3

@@ Abstract
In this paper, we consider a general observation model for restless multi-armed bandit problems. The operation of the player needs to be based on certain feedback mechanism that is error-prone due to resource constraints or environmental or intrinsic noises. By establishing a general probabilistic model for dynamics of feedback/observation, we formulate the problem as a restless bandit with a countable belief state space starting from an arbitrary initial belief (a priori information). We apply the achievable region method with partial conservation law (PCL) to the infinite-state problem and analyze its indexability and priority index (Whittle index). Finally, we propose an approximation process to transform the problem into which the AG algorithm of Niño-Mora and Bertsimas for finite-state problems can be applied to. Numerical experiments show that our algorithm has an excellent performance.
@@
2Jiangsu National Center for Applied Mathematics, Nanjing, 210093, Jiangsu, China.
3Bank of Communications, Shanghai, 200120, China.
@@
Keywords Restless bandit  $\cdot$ POMDP  $\cdot$ Countable state space  $\cdot$ Partial conservation law  $\cdot$ Whittle index
@@ 1 Introduction
Multi-armed bandit (MAB) is a classic operations research problem that involves a learner making choices among actions with uncertain/random rewards in order to maximize the expected return in long-run. The MAB problem is often associated with the important exploration-exploitation tradeoff and was initially proposed by Robbins (1952). Starting from the classic stochastic scheduling problem, the ongoing development of various MAB models makes it applicable to a wide range of practical fields, including clinical trials, recommendation systems, cognitive communications, and financial investments (Gittins, 1979; Berry and Fristedt, 1985; Press, 2009; Farias and Madan, 2011; Hoffman et al., 2011; Shen et al., 2015).
@@
In the classic Bayesian MAB problem, the bandit machine has a total of $N$ arms and one player pulling an arm in each time slot. After a player selected one of these arms and activated it, a random reward will be accrued depending on the activated arm and its current state. In this process, all states are visible, and only the state of the arm being activated changes according to a Markov chain. The goal of the MAB problem is to search a policy that maximizes the long-term cumulative reward. Gittins (1979) proved that the classic MAB problem can be solved optimally by an index policy, referred to as Gittins index, and the player only needs to activate the arm with the largest index at each moment. The index policy significantly reduces the complexity of the problem from being exponential with the number of arms to being linear. After that, Whittle (1988) extended the MAB problem to the restless MAB (RMAB) model, in which $K$ arms can be activated at each moment, and the arms not chosen may also undergo state transitions over time. However, for the general RMAB problem with a finite state space, Papadimitriou and Tsitsiklis (1994) showed that the computational complexity of finding the optimal policy is already PSPACE-hard. Fortunately, Whittle (1988) generalized Gittins index to Whittle index that provides a possible solution to RMAB by considering Lagrangian relaxation and duality. Accordingly, Whittle index is optimal under a relaxed constraint that the time-average number of activated arms is $K$. For the original problem, the Whittle index policy has been proven to be asymptotically optimal per-arm wise as the number of arms goes to infinity under certain conditions (Weber and Weiss, 1990a, b). Nonetheless, not every RMAB is indexable, i.e., Whittle index may not exist. And the establishment of indexability is itself a difficult problem. Even if indexability is proven for a particular RMAB, analytical solutions of the Whittle index function can still be hard to obtain. For some important categories of RMAB models, Whittle indexability and its strong performance have been demonstrated in the literature, such as the dual-speed bandit problem and partially observable RMAB (Glazebrook et al., 2002; Ahmad et al., 2009; Liu and Zhao, 2010; Gittins and Weber, 2011; Liu, 2025; Liu et al., 2024).
@@
In this paper, we extend the partially observable RMAB considered in Liu et al. (2010, 2024) to general observation models. The previous work only considers special classes of observation errors and noises within the framework of partially observation Markov decision processes (POMDP) (Zhao and Sadler, 2007; Liu et al., 2010, 2024). But general algorithms for POMDP often suffer from curse of dimensionality and becomes cumbersome when the value functions for dynamic programming are too complex (Smallwood and Sondik, 1973; Sondik, 1978). In order to deal with general observation models, we need an alternative methodology to design efficient algorithms for systems with large sets of parameters. In the interdisciplinary field of operations research and stochastic optimization, an analytical method called “Achievable Region” emerged in the 1990s. Interestingly, this method transforms a time-series optimization problem into linear programming (LP) problems that can be solved efficiently (Coffman Jr and Mitrani, 1980; Federgruen and Groenevelt, 1988a; Federgruen and Groenevelt, 1988b; Gelenbe and Mitrani, 2010). The challenge of using this method is mainly on proving the feasibility of such transformation (Bertsimas, 1995). Referred to as the general conservation law (GCL), Bertsimas and Niño-Mora (1996) proposed the required structure of performance measures for the achievable region to be a polyhedron called extended polymatroid. Based on GCL, they extended Klimov’s algorithm to several stochastic scheduling problems, including the classic MAB problem (Klimov, 1975; Division and Tsoucas, 1991). By relaxing the GCL restrictions, Niño-Mora (2001) proposed the partial conservation law (PCL) and specified conditions for an RMAB to satisfy PCL and subsequently lead to the numerical calculation of Whittle index. Later on, Niño-Mora (2002) offered an economic explanation of PCL-indexability and its relation to Whittle indexability where Whittle index can be interpreted as the optimal marginal cost rate. A more comprehensive summary of this work can be found in Niño-Mora (2007). However, all such work considers finite state spaces except that Frostig and Weiss (1999) extended the GCL framework for classic MAB to countable state spaces. In this paper, we extend the PCL framework to the RMAB with a general parametrization of observation model and an infinite state space. The main challenge of this extension is due to the high-dimensional probability state space (belief states) with complex dynamics for transitions over time. By establishing the weak duality between LP formulations, we build the PCL framework for analyzing our RMAB and subsequently design an efficient algorithm for calculating Whittle index when PCL-indexability is satisfied. Finally, we demonstrate the superior performance of our algorithm by numerical simulations.
@@ 2.1 Model Formulation
In this section, we begin with the formulation of the RMAB problem with general observation models. Assume the system has $N$ arms and the state space of the $n$th arm is $\mathcal{M}_{n}$. Therefore the entire system state space is $\mathcal{M}=\prod_{n=1}^{N}\mathcal{M}_{n}$. At each moment, the actual state of each arm undergos a state transition based on its own Markov probability transition matrix, and we will select $K$ arms to activate. For those activated arms, we can observe their states (with errors) and accrue some reward dependent on the observed states and the actual states. For those arms that are not activated, we can neither observe their states nor obtain reward from them, and their states still transit over time. Let $S_{n}(t),O_{n}(t)\in\mathbb{Z}^{+}$ be respectively the true and observed states of arm $n$ in slot $t$ and $A(t)$ the set of activated arms in slot $t$. First consider a single-arm process (i.e. $N=1$ and $\mathcal{M}=\mathcal{M}_{n}$) and drop the subscript $n$ (sometimes the time index $t$ is also dropped if referring to the same time index) for convenience. Suppose the state transition matrix, error matrix and reward matrix are $P=\{p_{ij}\}_{\mathcal{M}\times\mathcal{M}}$, $E=\{\varepsilon_{ij}\}_{\mathcal{M}\times\mathcal{M}}$ and $R=\{r_{ij}\}_{\mathcal{M}\times\mathcal{M}}$ respectively. For error matrix, $\varepsilon_{ij}$ represents the probability that the observed state is $j$ when the true state is $i$, that is, $\varepsilon_{ij}=P(O=j|S=i)$, while $r_{ij}$ represents the reward obtained when the true state is $i$ and the observed state is $j$ (different observations may lead to different sub-actions and thus different rewards following). Denoted by ${\Omega}_{a}=\{(\omega_{1},\cdots,\omega_{M})|\sum_{i=1}^{M}\omega_{i}=1,\omega_{i}\geq 0,1\leq i\leq M:=|\mathcal{M}|\}$ the belief space, for each $\boldsymbol{\omega}(t)=(\omega_{1}(t),\cdots,\omega_{M}(t))\in{\Omega_{a}}$, $\omega_{i}(t)$ represents the conditional probability that the true state is $i$ in slot $t$ (based on past observations). If the current belief state is $(\omega_{1},\cdots,\omega_{M})$, the current expected reward is $\sum_{i=1}^{M}\omega_{i}(\sum_{j=1}^{M}\varepsilon_{ij}r_{ij})$. In addition, we may sometimes receive additional feedback related to the true state and the immediate reward obtained that helps us better estimate the actual arm state (e.g., ACK/NAK in a communication channel to indicate whether the data was successfully transmitted (Liu et al., 2024)). Assume that there are $L$ feedback states (positive integers) in total that encompass all possible observations, and denote the feedback state at time $t$ by $F(t)~{}(L\leq M^{2})$. Define $\rho_{ij}=P(F(t)=j|S(t)=i)$. By Bayes rule, we have
@@
In a simpler scenario where there is no additional feedback but only the observed states, the update rule becomes
@@
In the extreme case when we cannot observe any state after activating an arm, we simply treat the obtained reward (might be 00) as the feedback. By normalizing the immediate reward $r(t)$ to take only positive integer values, we have
@@
For the restless multi-armed bandit model, the goal is to find a policy $\pi$ which maps the belief states of all arms into an active set $A(t)$ in slot $t$ that maximize the long-term expected discounted reward. In other words, if we denote the reward obtained by the $n$-th arm in slot $t$ by $r_{n}(t)$, then our objective is
@@
where $10\leq\beta<$ 1 is the discount factor and $\boldsymbol{\omega}_{n}(t)$ is the belief state (vector) of arm $n$ in slot $t$. In this problem, the diversity of states, choices, errors makes the problem highly complex. In RMAB problems, searching for an easily computable priority index policy is the mainstream. The core idea is to assign an index (a real number) to the current state of each arm, and then activate those arms with top $K$ large indexes. The goal of this paper is to theoretically characterize the conditions when such an index policy exists and provide a detailed algorithm for efficiently computing the index function (if it exists).
@@ 2.2 Whittle Index
Whittle (1988) relaxed the constraint on the exact number of arms activated in each slot, requiring only that the expected number of arms activated per slot on average (in the sense of discounted time) is $K$, i.e.
@@
where $\boldsymbol{\Omega}(t)=(\boldsymbol{\omega}_{1}(t),\cdots,\boldsymbol{\omega}_{N}(t))$ with $\boldsymbol{\omega}_{n}(t)$ as the belief state (vector) for arm $n$ at time $t$. Thus the Lagrange optimization problem can be written as
@@
The above problem can be decomposed into $N$ independent subproblems, that is, for any $1\leq n\leq N$,
@@
The explanation of the optimization problem above is that for each arm, when it is not selected (made passive), we will receive a subsidy $\lambda$. Since the problems above are independent, we just need to consider the single-arm case. For notation simplicity, we drop the arm index $n$ from now on. For a given arm, the optimal policy for the relaxed optimization problem divides the arm state space into two subset (here the arm state space is the arm belief state space): active set $\mathcal{A}(\lambda)$ and passive set $\mathcal{P}(\lambda)$. Specifically, $\mathcal{P}(\lambda)$ contains all belief states in which the optimal choice is passive when the subsidy is $\lambda$. In particular, for a certain state $s$, if both active and passive actiosn are optimal, we include it in $\mathcal{P}(\lambda)$, and $\mathcal{A}(\lambda)$ is just the complement of $\mathcal{P}(\lambda)$ in the entire state space. Under the concept of passive set, Whittle indexability can be stated as follows:
@@
An arm is indexable if the passive set $\mathcal{P}(\lambda)$ increases from $\emptyset$ to the whole state space as the subsidy $\lambda$ increases from $\infty$- ∞ to +$\infty$+ ∞. The RMAB problem is indexable if every decoupled problem is indexable.

### v4

# General Formulation and PCL-Analysis for Restless Bandits with Limited Observability

@@ Abstract
In this paper, we consider a general observation model for restless multi-armed bandit problems. The operation of the player is based on the past observation history that is limited (partial) and error-prone due to resource constraints or environmental or intrinsic noises. By establishing a general probabilistic model for dynamics of the observation process, we formulate the problem as a restless bandit with an infinite high-dimensional belief state space. We apply the achievable region method with partial conservation law (PCL) to the infinite-state problem and analyze its indexability and priority index (Whittle index). Finally, we propose an approximation process to transform the problem into which the AG algorithm of Niño-Mora (2001) for finite-state problems can be applied. Numerical experiments show that our algorithm has excellent performance.
@@
Keywords: Restless bandit; POMDP; Countable state space; Partial conservation law; Whittle index
@@ 1 Introduction
Multi-armed bandit (MAB) is a classic operations research problem that involves a learner making choices among actions with uncertain/random rewards in order to maximize the expected return in long-run. The MAB problem is often associated with the important exploration-exploitation tradeoff and was initially proposed by robbins1952some. Starting from the classic stochastic scheduling problem, the ongoing development of various MAB models makes it applicable to a wide range of practical fields, including clinical trials, recommendation systems, cognitive communications, and financial investments (gittins1979bandit; berry1985bandit; press2009bandit; farias2011irrevocable; hoffman2011portfolio; shen2015portfolio).
@@
In the classic Bayesian MAB problem, the bandit machine has a total of $N$ arms and one player pulling an arm in each decision epoch. After a player selected one of these arms and activated it, a random reward is accrued depending on the activated arm and its current state. In this process, all states are visible, and only the state of the arm being activated changes according to a Markov chain. The goal of the MAB problem is to search a policy that maximizes the long-term cumulative reward. gittins1979bandit proved that the classic MAB problem can be solved optimally by an index policy, referred to as Gittins index, and the player only needs to activate the arm with the largest index at each moment. The index policy significantly reduces the complexity of the problem from being exponential with the number of arms to being linear. After that, whittle1988restless extended the MAB problem to the restless MAB (RMAB) model, in which $K$ arms can be activated at each moment, and the arms not chosen may also undergo state transitions over time. However, for the general RMAB problem with a finite state space, papadimitriou1994complexity showed that the computational complexity of finding the optimal policy is already PSPACE-hard. Fortunately, whittle1988restless generalized Gittins index to Whittle index that provides a possible solution to RMAB by considering Lagrangian relaxation and duality. Accordingly, Whittle index is optimal under a relaxed constraint that the time-average number of activated arms is $K$. For the original problem, the Whittle index policy has been proven to be asymptotically optimal per-arm wise as the number of arms goes to infinity under certain conditions (weber1990index; weber1991index). Nonetheless, not every RMAB is indexable, i.e., Whittle index may not exist. And the establishment of indexability is itself a difficult problem. Even if indexability is proven for a particular RMAB, analytical solutions of the Whittle index function can still be hard to obtain. For some important categories of RMAB models, Whittle indexability and its strong performance have been demonstrated in the literature, such as the dual-speed bandit problem and partially observable RMAB (glazebrook2002index; ahmad2009optimality; liu2010indexability; gittins2011multi; larranaga2016rmab; AyestaEtal2021MMOR; LiuEtal24MMOR; fu2025rmab; liu2021index).
@@
In this paper, we extend the partially observable RMAB considered in liu2010dynamic; LiuEtal24MMOR; liu2021index to general observation models. The previous work only considers special classes of observation errors and noises within the framework of partially observable Markov decision processes (POMDP) (zhao2007survey; liu2010dynamic; LiuEtal24MMOR). But general algorithms for POMDP often suffer from curse of dimensionality and become cumbersome when the value functions for dynamic programming are too complex to be precisely solved or well approximated (smallwood1973optimal; sondik1978optimal). For special cases, the state transition dynamics yield a simple structure (e.g., threshold type) of the optimal policy such that the value functions can be efficiently evaluated or even solved in closed-form (gittins2011multi; liu2010dynamic; LiuEtal24MMOR; liu2021index). In order to deal with general observation models, we need an alternative methodology to design efficient algorithms for systems with large sets of parameters. In the interdisciplinary field of operations research and stochastic optimization, an analytical method called “Achievable Region” emerged in the 1990s. Interestingly, this method transforms a time-series optimization problem into linear programming (LP) problems that can be solved efficiently (coffman1980characterization; federgruen1988characterization; federgruen1988m; gelenbe2010analysis). The challenge of using this method is mainly on proving the feasibility of such transformation (bertsimas1995achievable). Referred to as the general conservation law (GCL), bertsimas1996conservation proposed the required structure of performance measures for the achievable region to be a polyhedron called extended polymatroid. Based on GCL, they extended Klimov’s algorithm to several stochastic scheduling problems, including the classic MAB problem (klimov1975time; thomas1991region). By relaxing the GCL restrictions, nino2001restless proposed the partial conservation law (PCL) and specified conditions for an RMAB to satisfy PCL and subsequently lead to the numerical calculation of Whittle index. Later on, nino2002dynamic offered an economic explanation of PCL-indexability and its relation to Whittle indexability where Whittle index can be interpreted as the optimal marginal cost rate. A more comprehensive summary of this work can be found in nino2007dynamic. However, all such work considers finite state spaces except for frostig1999four who extended the GCL framework for classic MAB to countable state spaces. As detailed in nino2001restless, the PCL framework is a significant generalization of GCL and works for a much broader class of Whittle-indexable RMAB compared to the GCL. Therefore, the complexity in building the PCL framework is much higher than that of GCL, as well as its extension to infinite state spaces. For example, GCL has a strong requirement that the full state combination set must satisfy a conservation law for the existence and optimality of the index policy; while PCL only requires a state combination subset satisfying a generalized conservation law. This significantly broadened the applicability of LP approach for solving bandit problems. Nevertheless, why the relaxed condition by PCL still ensures an optimal index policy is a deep theorem (nino2001restless), especially in the case of infinite states while state subsets become too diverse and also infinite.
The main contribution of this paper is to establish the validity of PCL-framework for the general infinite-state RMAB model by carefully dealing with the limiting scenarios and fully exploring the intrinsic rich structure in the analysis of this high-dimensional infinite-state LP problem. The main challenge of this extension is due to the high-dimensional probability state space (belief states) with complex dynamics for transitions over time. By establishing the weak duality between LP formulations, we build the PCL framework for analyzing our RMAB and subsequently design an efficient algorithm for calculating Whittle index when PCL-indexability is satisfied. Finally, we demonstrate the superior performance of our algorithm by numerical simulations. We point out that nino2006restless and nino2020mor also considered PCL for special instances of RMAB with infinite state spaces. Specifically, nino2006restless formulated the $M/G/1$ queuing control problem as an RMAB with a countable state space equipped with a simple state transition rule and a convex holding cost function where the family of threshold policies was focused on to analyze the indexability. nino2020mor extended the work of liu2010dynamic for partially but error-free observable RMAB to the real-interval state space, i.e., the state space has dimension $1$. Furthermore, only the family of threshold policies was considered but rather than proving the optimality of threshold policies before establishing indexability as in liu2010dynamic, nino2020mor gave sufficient conditions to obtain both the optimality of threshold policies and indexability at one time. AyestaEtal2021MMOR adopted the time-average objective function with perfect observations. Under the ergodicity condition, the authors utilized the state transition dynamics to propose an improved exhaustive-search algorithm to check indexability and compute Whittle index. However, the algorithm works only for the case where the threshold policy is optimal in the single-armed problem with subsidy. Furthermore, the complexity of the algorithm is very high due to the exhaustive search in state combinations and numerical implementation of the algorithm still requires an approximate finite state space. In this paper, we consider the high-dimensional RMAB without any constraint on state dimension or transition or policy class and use linear programming to efficiently finish indexability verification and computation, as the first modeling and PCL analysis of RMAB with general observability.
Table 1: Main Notations
NN number of arms
$n$ arm index
$\mathcal{M}_{n}$ arm state space
$\mathcal{M}$ joint state space of all arms
$K$ number of arms to select at each decision epoch
$t$ decision epoch index
$S$ state of an arm
$O$ observed state of an arm
$p_{ij}$ transition probability between arm states
$\epsilon_{ij}$ probability of observed state to be $j$ given the true state being $i$
$r_{ij}$ immediate reward when the observed state is $j$ and the true state is $i$
$\omega_{i}$ probability the arm is in state $i$ given the past observations
$F$ feedback state
$\rho_{ij}$ probability of getting immediate reward $j$ given the true state is $i$
$\beta$ discount factor
$\lambda$ subsidy for passivity in the single-arm problem
$\mathcal{P}(\lambda)$ passive set under subsidy $\lambda$
$W(\cdot)$ Whittle index function
$\mathcal{B}$ belief update operator
$p_{ij}^{k}$ probability that belief state transits to $\boldsymbol{\omega}_{j}$ given $\boldsymbol{\omega}_{i}$ and action $a=k$ ($k=0$ for passive and $k=1$ for active)
$x_{\boldsymbol{\omega}}^{a}(\pi,\boldsymbol{\omega_{i}})$ expected total discounted time that $\boldsymbol{\omega}$ is applied with action $a$ under policy $\pi$ and initial belief $\boldsymbol{\omega_{i}}$
$\Omega(\boldsymbol{\omega_{0}})$ set of all possible belief states given initial belief $\boldsymbol{\omega_{0}}$
$\tilde{\Omega}$ feasible set family
$\Omega_{\boldsymbol{\omega_{i}}}^{C}$ an element (belief state subset) of full subset chain $C$ indexed by $\boldsymbol{\omega_{i}}$
@@ 2.1 Model Formulation
In this section, we begin with the formulation of the RMAB problem with general observation models. The main notations adopted in this paper can be found in Table 1. Assume the system has $N$ arms and the state space of the $n$-th arm is $\mathcal{M}_{n}$. Therefore the joint state space of all arms is $\mathcal{M}=\prod_{n=1}^{N}\mathcal{M}_{n}$. At each moment, the actual state of each arm undergoes a state transition based on its own Markov probability transition matrix, and we will select $K$ arms to activate. For those activated arms, we can observe their states (with errors) and accrue some reward dependent on the observed states and the actual states. For those arms that are not activated, we can neither observe their states nor obtain reward from them, and their states still transit over time. This assumption was motivated by practical scenarios where not selected arm does not provide reward or incur cost. For example, in the opportunistic spectrum access problem, if the user does not select an arm (channel) to sense and transmit data, then no observation or reward is received from that channel (see below and Sec. 3.1 for details). Relaxing this assumption will be considered in the future work. Let $S_{n}(t),O_{n}(t)\in\mathbb{Z}^{+}$ be respectively the true and observed states of arm $n$ in decision epoch $t$ and $A(t)$ the set of activated arms in $t$. First consider a single-arm process (i.e. $N=1$ and $\mathcal{M}=\mathcal{M}_{n}$) and drop the subscript $n$ (sometimes the time index $t$ is also dropped if referring to the same time index) for convenience. Suppose the state transition matrix, error matrix and reward matrix are $P=\{p_{ij}\}_{\mathcal{M}\times\mathcal{M}}$, $E=\{\varepsilon_{ij}\}_{\mathcal{M}\times\mathcal{M}}$ and $R=\{r_{ij}\}_{\mathcal{M}\times\mathcal{M}}$ respectively. For error matrix, $\varepsilon_{ij}$ represents the probability that the observed state is $j$ when the true state is $i$, that is, $\varepsilon_{ij}=P(O=j|S=i)$, while $r_{ij}$ represents the reward obtained when the true state is $i$ and the observed state is $j$ (different observations may lead to different sub-actions and thus different rewards following). Denote by ${\Omega}_{a}=\{(\omega_{1},\cdots,\omega_{M})|\sum_{i=1}^{M}\omega_{i}=1,\omega_{i}\geq 0,1\leq i\leq M:=|\mathcal{M}|\}$ the belief space, for each $\boldsymbol{\omega}(t)=(\omega_{1}(t),\cdots,\omega_{M}(t))\in{\Omega_{a}}$, $\omega_{i}(t)$ represents the conditional probability that the true state is $i$ in epoch $t$ (based on past observations). If the current belief state is $(\omega_{1},\cdots,\omega_{M})$, the current expected reward is $\sum_{i=1}^{M}\omega_{i}(\sum_{j=1}^{M}\varepsilon_{ij}r_{ij})$. In addition, we may sometimes receive additional feedback related to the true and observed states that helps us better estimate the actual arm state (e.g., ACK/NAK in a communication channel to indicate whether the data was successfully transmitted (LiuEtal24MMOR)). Assume that there are $L$ feedback states (positive integers dependent only on the true state and the observed state) in total that encompass all possible observations, and denote the feedback state at time $t$ by $F(t)~(L\leq M^{2})$. For example, in the opportunistic spectrum access (OSA) problem, the user first chooses an arm (channel) to sense. If the channel is observed to be available ($O(t)=1$), then the user sends a data packet to the receiver. If the channel is truly available ($S(t)=1$), then the receiver successfully receives this data packet and sends back an acknowledgment ($F(t)=1$) in the end of this decision epoch. In this case, the user knows the channel state ($S(t)=1$). However, due to sensor errors, the user may sense the channel and find it to be unavailable ($O(t)=0$) while the channel is actually available ($S(t)=1$). Then the user will not send any data packet and certainly no acknowledgment will be obtained ($F(t)=0$). In this case, the user never knows the true state of the channel in this decision epoch. By Bayes rule, we have
@@
In a simpler scenario where there is no additional feedback but only the observed states (e.g., no acknowledgment is sent back in the OSA problem), the update rule becomes
@@
In the extreme case when we cannot observe any state after activating an arm, we simply treat the obtained reward (might be 0) as the feedback. In general, the feedback state informs the player about the immediate reward obtained. By normalizing the immediate reward $r(t)$ to take only positive integer values ($r_{ij}$’s are assumed to be rational for all i,ji,~j since $Q$ is dense in $R$ and provides arbitrary precision in practical scenarios, Theorem 1.20 (b) in rudin1976principles), we have
@@
For the restless multi-armed bandit model, the goal is to find a policy $\pi$ that maps the belief states of all arms into an active set $A(t)$ in epoch $t$ that maximize the long-term expected discounted reward. In other words, if we denote the reward obtained by the $n$-th arm in epoch $t$ by $r_{n}(t)$, then our objective is
@@
where $0\leq\beta<1$ is the discount factor and $\boldsymbol{\omega}_{n}(t)$ is the belief state (vector) of arm $n$ in epoch $t$. We point out that the initial belief state can be set arbitrarily but after it is given, the system state space is countable since the belief update happens only at discrete time indices. In this problem, the diversity of states, choices, errors makes the problem highly complex. In RMAB problems, searching for an easily computable priority index policy is the mainstream. The core idea is to assign an index (a real number) to the current state of each arm, and then activate those arms with top $K$ large indexes. The goal of this paper is to theoretically characterize the conditions when such an index policy with near-optimal performance exists and provide a detailed algorithm for efficiently computing the index function (if it exists).
@@ 2.2 Whittle Index
whittle1988restless relaxed the constraint on the exact number of arms activated in each decision epoch, requiring only that the expected number of arms activated per epoch on average (in the sense of discounted time) is $K$, i.e.
@@
where $\boldsymbol{\Omega}(t)=(\boldsymbol{\omega}_{1}(t),\cdots,\boldsymbol{\omega}_{N}(t))$ with $\boldsymbol{\omega}_{n}(t)$ as the belief state (vector) for arm $n$ at time $t$. Thus the Lagrange relaxation of (10) can be written as
@@
Note that only the class of admissible (measurable) policies are considered, i.e., the action can only depend on past observations instead of future or unobserved information. The above problem can be decomposed into $N$ independent subproblems, that is, for any $1\leq n\leq N$,
@@
The explanation of the optimization problem above is that for each arm, when it is not selected (made passive), we will receive a subsidy $\lambda$. Since the problems above are independent, we just need to consider the single-arm case. In other words, the Lagrange relaxation decouples an $N$-dimensional problem into NN 11-dimensional problems, reducing the complexity with $N$ from exponential to linear! From this point on, we will focus on the single-arm problem with subsidy and drop the arm index $n$. For a given arm, the optimal policy for the relaxed optimization problem divides the arm state space into two subsets (here the arm state space is the arm belief state space): active set $\mathcal{A}(\lambda)$ and passive set $\mathcal{P}(\lambda)$. Specifically, $\mathcal{P}(\lambda)$ contains all belief states in which the optimal choice is passive when the subsidy is $\lambda$. In particular, for a certain state $s$, if both active and passive actions are optimal, we include it in $\mathcal{P}(\lambda)$, and $\mathcal{A}(\lambda)$ is just the complement of $\mathcal{P}(\lambda)$ in the entire state space. Under the concept of passive set, Whittle indexability can be stated as follows:
@@
An arm is indexable if the passive set $\mathcal{P}(\lambda)$ monotonically increases from $\emptyset$ to the whole state space as the subsidy $\lambda$ increases from -$\infty$ to $\infty$. The RMAB problem is indexable if every decoupled problem is indexable.